### v1

@@ A.1.4 Active AMPK, PIP3, AKT, and mTOR equations
A.1.5 Cytoplasmic and nuclear HIF1 equations
The hypoxia-inducible factor 1 (HIF1) gene is continuously transcribed and translated, yet its expression levels remain undetectable under normoxia due to its rapid degradation via the ubiquitin-proteasome pathway in an oxygen-dependent process [115, 75, 35]. However, with the activation of the growth factor signalling pathway, mTOR enhances HIF1 protein levels through activation of cap-dependent translation and further facilitates its nuclear translocation [26, 115, 75, 130, 109]. Let and denote the concentrations of cytoplasmic and nuclear HIF1, respectively.
@@ A.1.5 Cytoplasmic and nuclear HIF1 equations
A.1.6 GLUT1, GLUT3, PDK1/3, PDK2, and LDH equations
Glucose transporters (GLUT1 and GLUT3), along with pyruvate dehydrogenase kinases (PDK1, PDK2, and PDK3) and lactate dehydrogenase (LDH), play crucial roles in regulating glucose metabolism pathways. Under normal conditions, these proteins are produced at baseline rates, but their production can be significantly influenced by the activation of HIF1 or p53, which can either stimulate or inhibit their synthesis. Specifically, GLUT1 and GLUT3 are vital for cellular glucose uptake [99, 106, 1, 83], a process enhanced by HIF1 but dampened by p53 [1, 99, 56]. PDK kinases, which interfere with the conversion of pyruvate to acetyl-CoA in mitochondria [120, 119, 92], are differentially regulated: HIF1 promotes the expression of PDK1 and PDK3 [4, 68, 120, 62], disrupting pyruvate oxidation, while p53 acts to diminish PDK2 activity [4, 66]. LDH enzyme, pivotal for converting pyruvate to lactate in the cytoplasm, receives support from HIF1, facilitating the anaerobic glycolysis pathway [115]. This intricate gene regulation network highlights the dynamic interplay between HIF1 and p53 in modulating glucose metabolic pathways. Let the concentration of cytoplasmic GLUT1, cytoplasmic GLUT3, mitochondrial PDK1 and 3, mitochondrial PDK2, and cytoplasmic LDH be denoted by , , , , and , respectively.
@@ A.1.6 GLUT1, GLUT3, PDK1/3, PDK2, and LDH equations
A.1.7 Active and inactive PDH equations
Pyruvate dehydrogenase (PDH) is an enzyme that facilitates the first step towards glucose respiration by catalysing pyruvate oxidative decarboxylation into acetyl-CoA within the mitochondria [120, 119, 92]. However, this enzyme’s function is negatively regulated by the PDK family, which phosphorate PDH complex at three different sites, inhibiting its activity [120, 119, 92]. Let the concentration of active and inactive mitochondrial PDH be denoted by and , respectively.
@@ A.1.7 Active and inactive PDH equations
A.1.8 Metabolic equations
Glucose metabolism involves three main stages: starting with glycolysis, advancing to the tricarboxylic acid (TCA) cycle, and finishing with the electron transport chain (ETC).
[…]
Glucose is transported into and out of the cells by specific glucose transporters located in the cell membrane [99, 106, 1, 83], namely GLUT1 and GLUT3 in our model. Once inside the cell, glucose undergoes an irreversible conversion to glucose-6-phosphate (G6P), consuming one ATP molecule [35]. Then the G6P can either proceed through glycolysis or enter the pentose phosphate pathway (PPP) [53], a decision influenced by TIGAR which inhibits the enzyme catalysing the third step in the glycolysis pathway, diminishing glycolysis flux [13, 67]. Continuing with glycolysis converts a molecule of G6P into two molecules of pyruvate, yielding three net ATP and two reduced nicotinamide adenine dinucleotide (NADH) molecules [115, 35]. Following this, the produced pyruvate can either be converted into lactate in the cytoplasm by LDH enzyme, oxidising NADH back to NAD+, or transported into the mitochondria in the presence of oxygen for further energy production through oxidative phosphorylation (OXPHOS) [115, 35].
[…]
Within the mitochondria, the PDH complex irreversibly catalyses the oxidative decarboxylation of pyruvate to acetyl-CoA, concurrently reducing one NAD+ into NADH [120, 119, 92]. Acetyl-CoA subsequently enters the TCA cycle, undergoing a series of chemical reactions that generate one ATP, three NADH, and one reduced flavin adenine dinucleotide (FADH2) molecule [87].
[…]
In the last stage, NADH and FADH2 are oxidised back into NAD+ and FAD via protein complexes in the inner mitochondrial membrane [7]. This oxidation process involves electron transfer from NADH and FADH2 across these complexes, during which protons are pumped from the mitochondrial matrix to the intermembrane space. This action creates a proton gradient that drives protons back into the matrix, facilitating ATP production [7]. Given that NADH and FADH2 contribute to the pumping of approximately ten and six protons, respectively, and the movement of every four protons generates one ATP molecule, our model estimates that NADH and FADH2 respectively produce 2.5 and 1.5 ATP molecules [7]. Among these complexes, Complex IV acts as the final electron acceptor, channelling the electrons to molecular oxygen by using 0.5 oxygen (O2) molecules for each pair of electrons received from NADH or FADH2 [7]. However, the activity of Complex IV is regulated by SCO2, a p53-regulated gene essential for its proper assembly and maturation. Thus, deficiency of SCO2 can impair Complex IV functionality and disrupt the electron flow within the ETC [84, 118].
@@ A.1.8 Metabolic equations
A.2 Discussion of Parameter Values
@@
A.3 Parameter Robustness Analysis
To investigate our model robustness against parameter uncertainties, we conducted a sensitivity analysis (), assessing how changes in each parameter () influence key system outcomes () across both cancer cell phenotypes. During this analysis, each parameter was individually adjusted by 10% above or below its baseline value, while all other parameters remained constant, as detailed in Table 11.
The data indicate that cancer cells with intact p53 are more sensitive to parameter variations, especially in terms of glucose consumption and lactate production, when compared to p53-mutant cells. Conversely, processes like oxygen consumption and ATP production were more robust, showing no sensitivity. The parameters significantly affecting these responses were primarily associated with the dynamics of p53 activation and deactivation by WIP1 (, , , and ), and those regulating growth factor signalling pathways (, , , , , , , and ). Notably, the oxygen-dependent degradation rate of HIF1 () and the rate of glucose transport () were critical for both cell types.
| Cancer cells (p53) | Cancer cells (p53) | ||||||||||||||||
| Glucose | Lactate | Oxygen | ATP | Glucose | Lactate | Oxygen | ATP | ||||||||||
| Parameter | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | |
| k | -0.14 | -0.17 | -0.16 | -0.20 | 0.01 | 0.02 | -0.04 | -0.04 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -1.14 | -1.00 | -1.32 | -1.16 | 0.12 | 0.12 | -0.29 | -0.24 | - | - | - | - | - | - | - | - | |
| k | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.84 | 1.13 | 0.98 | 1.30 | -0.10 | -0.12 | 0.20 | 0.28 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.12 | 0.13 | 0.14 | 0.15 | -0.01 | -0.01 | 0.03 | 0.03 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.13 | 0.14 | 0.16 | 0.16 | -0.01 | -0.01 | 0.03 | 0.03 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.03 | 0.03 | 0.03 | 0.03 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.11 | 0.11 | 0.13 | 0.13 | -0.01 | -0.01 | 0.03 | 0.03 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.01 | 0.01 | 0.01 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -0.13 | -0.15 | -0.15 | -0.17 | 0.01 | 0.02 | -0.03 | -0.04 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.06 | 0.06 | 0.07 | 0.07 | -0.01 | -0.01 | 0.01 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.80 | 1.06 | 0.92 | 1.23 | -0.09 | -0.11 | 0.19 | 0.26 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -1.02 | -0.92 | -1.17 | -1.06 | 0.10 | 0.11 | -0.25 | -0.22 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -0.50 | -0.46 | -0.55 | -0.50 | 0.00 | 0.00 | -0.15 | -0.13 | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -1.09 | -0.92 | -1.20 | -0.98 | 0.02 | 0.00 | -0.31 | -0.26 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 1.00 | 1.39 | 1.05 | 1.53 | 0.00 | -0.04 | 0.28 | 0.39 | 0.00 | 0.01 | 0.00 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | -0.03 | -0.03 | 0.11 | 0.12 | 0.07 | 0.08 | 0.00 | 0.00 | -0.04 | -0.04 | 0.32 | 0.37 | 0.16 | 0.19 | |
| k | 0.00 | 0.00 | -0.03 | -0.03 | 0.14 | 0.15 | 0.09 | 0.10 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | 0.06 | 0.06 | -0.25 | -0.27 | -0.16 | -0.17 | 0.00 | 0.00 | 0.04 | 0.04 | -0.33 | -0.35 | -0.17 | -0.18 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | -0.06 | -0.06 | -0.01 | -0.01 | -0.03 | -0.03 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | 0.05 | 0.06 | 0.01 | 0.01 | 0.03 | 0.04 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -0.01 | -0.02 | -0.02 | -0.02 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -0.01 | -0.01 | -0.01 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.02 | 0.02 | 0.02 | 0.02 | 0.00 | 0.00 | 0.01 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 1.00 | 1.39 | 1.05 | 1.53 | 0.00 | -0.04 | 0.28 | 0.39 | 0.00 | 0.01 | 0.00 | 0.01 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | -1.28 | -1.05 | -1.41 | -1.11 | 0.03 | 0.00 | -0.36 | -0.29 | -0.01 | 0.00 | -0.01 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.86 | 1.15 | 0.92 | 1.26 | 0.00 | -0.02 | 0.24 | 0.33 | 0.06 | 0.09 | 0.06 | 0.09 | 0.00 | 0.00 | 0.02 | 0.04 | |
| k | -1.04 | -0.94 | -1.14 | -0.99 | 0.02 | 0.00 | -0.30 | -0.26 | -0.08 | -0.06 | -0.08 | -0.07 | 0.00 | 0.00 | -0.03 | -0.03 | |
| k | 0.72 | 0.80 | 0.77 | 0.88 | 0.00 | -0.01 | 0.20 | 0.23 | 0.19 | 0.28 | 0.19 | 0.29 | -0.01 | -0.01 | 0.08 | 0.12 | |
| k | -0.73 | -0.79 | -0.79 | -0.84 | 0.01 | 0.00 | -0.21 | -0.22 | -0.25 | -0.20 | -0.26 | -0.21 | 0.01 | 0.01 | -0.10 | -0.08 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.34 | 0.34 | 0.37 | 0.37 | 0.00 | 0.00 | 0.10 | 0.10 | 0.26 | 0.26 | 0.27 | 0.27 | -0.01 | -0.01 | 0.10 | 0.11 | |
| k | 0.45 | 0.45 | 0.48 | 0.49 | 0.00 | 0.00 | 0.13 | 0.13 | 0.49 | 0.50 | 0.51 | 0.52 | -0.03 | -0.02 | 0.20 | 0.21 | |
| k | 0.10 | 0.12 | 0.11 | 0.13 | 0.00 | 0.00 | 0.03 | 0.04 | 0.07 | 0.08 | 0.07 | 0.09 | 0.00 | 0.00 | 0.03 | 0.03 | |
| k | -0.82 | -1.00 | -0.90 | -1.06 | 0.01 | 0.00 | -0.24 | -0.28 | -0.76 | -0.92 | -0.79 | -0.95 | 0.02 | 0.08 | -0.32 | -0.35 | |
| k | 0.11 | 0.11 | 0.12 | 0.12 | 0.03 | 0.03 | 0.05 | 0.05 | 0.13 | 0.13 | 0.13 | 0.13 | 0.01 | 0.02 | 0.06 | 0.07 | |
| k | 0.49 | 0.49 | 0.49 | 0.50 | 0.11 | 0.13 | 0.21 | 0.23 | 0.47 | 0.47 | 0.47 | 0.48 | 0.04 | 0.07 | 0.22 | 0.24 | |
| k | -0.55 | -0.67 | -0.57 | -0.68 | -0.15 | -0.14 | -0.26 | -0.28 | -0.54 | -0.66 | -0.56 | -0.67 | -0.08 | -0.05 | -0.28 | -0.31 | |
| k | 0.07 | 0.07 | 0.07 | 0.07 | 0.02 | 0.02 | 0.03 | 0.03 | 0.08 | 0.08 | 0.08 | 0.08 | 0.01 | 0.01 | 0.04 | 0.04 | |
| k | 0.30 | 0.30 | 0.31 | 0.31 | 0.07 | 0.08 | 0.13 | 0.14 | 0.29 | 0.29 | 0.29 | 0.30 | 0.03 | 0.04 | 0.14 | 0.15 | |
| k | -0.34 | -0.42 | -0.35 | -0.42 | -0.09 | -0.09 | -0.16 | -0.18 | -0.34 | -0.41 | -0.34 | -0.41 | -0.05 | -0.04 | -0.17 | -0.19 | |
| k | 0.00 | 0.00 | 0.01 | 0.00 | -0.02 | -0.02 | -0.02 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | -0.03 | -0.03 | -0.02 | -0.02 | |
| k | 0.00 | 0.00 | 0.01 | 0.00 | -0.02 | -0.02 | -0.02 | -0.01 | 0.00 | 0.00 | 0.01 | 0.00 | -0.05 | -0.04 | -0.02 | -0.02 | |
| k | 0.00 | 0.00 | -0.01 | -0.02 | 0.04 | 0.07 | 0.03 | 0.04 | 0.00 | 0.00 | -0.01 | -0.01 | 0.07 | 0.12 | 0.04 | 0.06 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | -0.01 | -0.01 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | -0.02 | -0.02 | -0.01 | -0.01 | |
| k | 0.00 | 0.00 | 0.02 | 0.02 | -0.07 | -0.08 | -0.04 | -0.05 | 0.00 | 0.00 | 0.00 | 0.00 | -0.02 | -0.03 | -0.01 | -0.01 | |
| k | 0.00 | 0.00 | 0.03 | 0.05 | -0.15 | -0.20 | -0.10 | -0.13 | 0.00 | 0.00 | 0.01 | 0.01 | -0.07 | -0.10 | -0.04 | -0.05 | |
| k | 0.00 | 0.00 | -0.06 | -0.05 | 0.27 | 0.22 | 0.17 | 0.14 | 0.00 | 0.00 | -0.01 | -0.01 | 0.12 | 0.10 | 0.06 | 0.05 | |
| k | 0.00 | 0.00 | -0.02 | -0.02 | 0.08 | 0.09 | 0.05 | 0.06 | 0.00 | 0.00 | -0.01 | -0.01 | 0.07 | 0.09 | 0.04 | 0.04 | |
| k | 0.00 | 0.00 | 0.01 | 0.01 | -0.06 | -0.05 | -0.04 | -0.03 | 0.00 | 0.00 | 0.01 | 0.01 | -0.11 | -0.08 | -0.05 | -0.04 | |
| k | 0.00 | 0.00 | -0.01 | -0.01 | 0.04 | 0.06 | 0.03 | 0.04 | 0.00 | 0.00 | -0.01 | -0.01 | 0.07 | 0.10 | 0.04 | 0.05 | |
[…]
| Cancer cells (p53) | Cancer cells (p53) | ||||||||||||||||
| Glucose | Lactate | Oxygen | ATP | Glucose | Lactate | Oxygen | ATP | ||||||||||
| Parameter | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | +10% | -10% | |
| k | 0.00 | 0.00 | 0.02 | 0.02 | -0.08 | -0.08 | -0.05 | -0.05 | 0.00 | 0.00 | 0.01 | 0.01 | -0.07 | -0.07 | -0.03 | -0.04 | |
| k | 0.98 | 0.98 | 0.98 | 1.02 | 0.19 | 0.28 | 0.39 | 0.48 | 0.96 | 0.97 | 0.97 | 1.00 | 0.05 | 0.18 | 0.43 | 0.52 | |
| k | 0.02 | 0.03 | 0.02 | 0.03 | 0.01 | 0.01 | 0.01 | 0.01 | 0.03 | 0.05 | 0.04 | 0.05 | 0.00 | 0.01 | 0.02 | 0.02 | |
| k | 0.00 | 0.00 | 0.12 | 0.15 | 0.02 | 0.03 | 0.07 | 0.09 | 0.00 | 0.00 | 0.06 | 0.08 | 0.00 | 0.00 | 0.04 | 0.05 | |
| k | 0.00 | 0.00 | -0.10 | -0.10 | -0.02 | -0.02 | -0.06 | -0.06 | 0.00 | 0.00 | -0.05 | -0.06 | 0.00 | 0.00 | -0.04 | -0.04 | |
| k | 0.00 | 0.00 | 0.06 | 0.07 | -0.27 | -0.30 | -0.18 | -0.19 | 0.00 | 0.00 | 0.03 | 0.03 | -0.24 | -0.27 | -0.13 | -0.14 | |
| k | 0.00 | 0.00 | -0.05 | -0.05 | 0.21 | 0.23 | 0.13 | 0.15 | 0.00 | 0.00 | -0.02 | -0.02 | 0.20 | 0.22 | 0.10 | 0.11 | |
| k | 0.00 | 0.00 | -0.01 | -0.02 | 0.06 | 0.08 | 0.04 | 0.05 | 0.00 | 0.00 | -0.01 | -0.01 | 0.06 | 0.07 | 0.03 | 0.04 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | -0.05 | -0.06 | 0.23 | 0.26 | 0.14 | 0.16 | 0.00 | 0.00 | -0.02 | -0.02 | 0.18 | 0.20 | 0.09 | 0.10 | |
| k | 0.00 | 0.00 | -0.06 | -0.06 | 0.25 | 0.28 | 0.16 | 0.18 | 0.00 | 0.00 | -0.04 | -0.04 | 0.32 | 0.37 | 0.16 | 0.19 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | |
| k | 0.00 | 0.00 | 0.03 | 0.11 | 0.16 | 0.40 | 0.14 | 0.35 | -0.01 | -0.01 | 0.01 | 0.05 | 0.22 | 0.55 | 0.14 | 0.36 | |
[…]

### v2

@@ A.3.4 Active AMPK, PIP3, AKT, and mTOR equations
A.3.5 Cytoplasmic and nuclear HIF1 equations
Let and denote the concentrations of cytoplasmic and nuclear HIF1, respectively.
@@ A.3.5 Cytoplasmic and nuclear HIF1 equations
A.3.6 GLUT1, GLUT3, PDK1/3, PDK2, and LDH equations
Let the concentration of cytoplasmic GLUT1, cytoplasmic GLUT3, mitochondrial PDK1 and 3, mitochondrial PDK2, and cytoplasmic LDH be denoted by , , , , and , respectively.
@@ A.3.6 GLUT1, GLUT3, PDK1/3, PDK2, and LDH equations
A.3.7 Active and inactive PDH equations
Let the concentration of active and inactive mitochondrial PDH be denoted by and , respectively.
@@ A.3.7 Active and inactive PDH equations
A.3.8 Metabolic equations
@@
A.4 Discussion of Parameter Values
@@
A.5 Parameter Robustness Analysis
To investigate our model’s robustness against parameter uncertainties, we performed a global sensitivity analysis across both cancer cell phenotypes using Morris method [87], see Fig. 10 for p53+/+ and Fig. 11 for p53-/-.
Morris method evaluates the Elementary Effects (EEs) of input parameters by randomly sampling the input space and conducting a series of one-at-a-time (OAT) experiments. In each experiment, one input parameter is varied while the others are kept fixed, and the resultant change in the output is observed. This process is repeated multiple times to explore different regions of the input space.
Following these experiments, the mean () and standard deviation () of these EEs are determined, providing insights into the influence and interaction of each parameter on the model’s outcomes. The higher , the higher the overall impact of parameter on the output is, while a high standard deviation indicates that ’s impact on the output is variable depending on the region of the parameter space, suggesting non-linearity or interactions with other parameters.
Applying this method to the main model’s outcomes reveals that cancer cells with intact p53 are more sensitive to parameter variations, especially in terms of glucose consumption and lactate production, compared to p53-mutant cells. Conversely, processes like oxygen consumption and ATP production were more robust, with sensitivity indices below 0.5 in both cell types. This robustness suggests that these processes are maintained consistently to sustain cellular energy balance even under different conditions.
The parameters significantly affecting glucose consumption and lactate production outcomes in p53-wild type cells were primarily associated with the dynamics of p53 activation and deactivation by WIP1 (, , , and ), as well as those regulating growth factor signalling pathways (, , , , , and ). These pathways are intricately linked to the metabolic processes, and their sensitivity underscores the importance of p53 in maintaining metabolic homeostasis and responding to cellular stressors.
On the other hand, the oxygen-dependent degradation rate of HIF1 () and the rate of glucose transport () were critical for both cell types. These parameters, however, exhibited a low standard deviation, indicating that their impact is consistent and linear across different regions of the parameter space. This consistency implies that these parameters have a predictable and stable influence on cellular metabolism, regardless of the p53 status.